OpenStudy (anonymous):

need help with a 9th grade math assignment!!

OpenStudy (anonymous):

1. The graph: A coordinate plane with a colored in, closed figure. Each segment must be labeled. 2. The inequalities: Show the work for finding the inequality that defines each segment. 3. A point: Choose a point within the boundaries of the shape, and show algebraically that the point satisfies all of the inequalities.

OpenStudy (anonymous):

i already have my coordinates for the closed figure, which are (5,5) (11,5) (8,3) (8,7)

OpenStudy (sepeario):

And which part do you need help with?

OpenStudy (anonymous):

finding the inequalities

OpenStudy (sepeario):

@rvc

rvc (rvc):

do you have diagram for that ?

OpenStudy (anonymous):

no

rvc (rvc):

@Michele_Laino could you please help us.

OpenStudy (anonymous):

oops wait i think i understand it now, thanks anyways guys!~

rvc (rvc):

cool :)

OpenStudy (michele_laino):

for example, I consider the segment between points (5,5) and (11,5), so , that segment can be represented as a subset of the cartesian plane, as below: \[\Large \left\{ {\left( {x,y} \right)\left. {} \right|\;5 \leqslant x \leqslant 11,\quad y = 5} \right\}\]

OpenStudy (michele_laino):

other wise, we can apply this, more general formula: the segment between points P_1and P_2, can be written as below: \[\Large {P_1} + t\left( {{P_1} - {P_2}} \right)\] where t is a parameter, such that: \[\Large 0 \leqslant t \leqslant 1\] and the difference between 2 points, say P_!=(x_1,y_1) and P_2=(x_2,y_2), is defined as below: \[\Large {P_1} - {P_2} = \left( {{x_1} - {x_2},{y_1} - {y_2}} \right)\]

OpenStudy (michele_laino):

oops,..P_1, not P_!